OpenStudy (anonymous):

4x-2y=18 5x+3y=17 Solve to find the solution using the elimination method. Answer had to be in simplified fraction ordered pair

OpenStudy (smileyxl3):

x=4 and y=-7

OpenStudy (anonymous):

@agent0smith

hero (hero):

To solve using the elimination method: Multiply both sides of the the first equation by 3 Multiply both sides of the the second equation by 2 Then add both equations together and solve for x.

OpenStudy (agent0smith):

4x-2y=18 5x+3y=17 as Hero said, 3(4x-2y=18) +2(5x+3y=17) _______________ make sure you multiply EVERYTHING in the equation by those numbers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got it thxs